# SAM-I-Am: Semantic Boosting for Zero-shot Atomic-Scale Electron Micrograph Segmentation

### v1

@@ 1 Introduction
Unfortunately, it has been challenging to develop domain-specific analytics that extract meaningful information from raw microscope data (Kalinin et al., 2023b). A central problem in TEM analysis is the accurate and precise segmentation of microstructural image features (Akers et al., 2021; Stuckner et al., 2022). This task is difficult to automate because of the millions of potential crystal structures, defects, and artifacts in microscopy data.(Spurgeon et al., 2021) Complex interactions between the electron beam and sample (Spurgeon et al., 2021) resulting from varying imaging conditions, sample orientation, sample thickness, and detector configuration, greatly change the representation of an object in collected data. As a result, unknown images are often poorly described by (limited) prior datasets available.
This data sparsity is a fundamental barrier to the use of many ML-based computer vision approaches. While low-cost synthetic images generated via approaches such as multislice simulation (Rangel DaCosta et al., 2021)can be used to train models, these require extensive specific a priori assumptions about samples and imaging conditions. Alternatively, few-shot approaches based limited human-generated feature labels have shown some success but can fail to generalize (Akers et al., 2021). Thus, there is a critical need for new analytics to accelerate material discovery in TEM, particularly in emerging automated microscopy that combines analytics with theory-driven hypothesis generation (Liu et al., 2022).
Here we introduce an approach that harnesses the power of a foundation model for TEM image segmentation. Most state-of-the-art TEM-segmentation techniques use convolutional neural architectures such as U-Net(Ronneberger et al., 2015). Models are trained on microscopy data that is carefully labeled by domain experts with substantial time and cost. Moreover, it remains challenging for models trained in such manner to adapt to out-of-distribution data. In contrast, our approach does X.
@@ 1.1 Related work
(Akers et al., 2021) proposes a semi-supervised few-shot machine learning approach for TEM segmentation of three oxide material systems, (1) SrTiO / Ge epitaxial heterostructures, (2) LaSrFeO thin films, and (3) MoO nanoparticles. Given a fine-tuning set of less than 10 tiny crops (chips) of the image that represent micro-structural features, their model is able to make mask predictions comparable to those of domain experts. However, such masks are limited to chip-level representations. (Stuckner et al., 2022), proposes transfer learning from MicroNet (a large dataset of microscopy images) for improving the mean intersection over union (mIoU) of microscopy segmentation tasks. They show that models pre-trained on their dataset exhibited superior generalization to new microscopy images compared to models trained on other datasets like ImageNet.
@@ 2 Background
The remainder of the paper is structured as follows. Section 1.1 discusses the related works in TEM image segmentation. Section 3 discusses the overview and proposed system methodology. Section 4 presents various experiments and evaluation of the proposed method. Section 5 discusses our ablation study while section 6 summarizes the paper.
3 Overview
@@ 3.1 Limitations of SAM+
SAM+ provides a mechanism to specify parameters that control autoprompts. Behind the scenes, it dissects an input image into layers where each layer contains equal sized crops of the image obtained by applying a sliding crop window. Next, each crop belonging to layer is prompted by projecting (overlaying) an grid of points to identify potential regions. A typical parameter, like the ‘prompt grid length’, for instance, specifies the size of input prompt for the first layer (). Given prompt density for the first layer’s crop, crops of subsequent layers will be prompted with sparser points. i.e. .
Unfortunately, SAM+’s vanilla parameters are inadequate for achieving microstructure segmentation. Consider Fig. 1, which shows a sample with three distinct materials stacked horizontally. SAM+’s segmentation is on the left. For the middle layer (SrTiO), SAM+ generates two separate masks. However, the whole middle section is actually the same material ‘split’ by an antiphase boundary defect (typically of special interest). This shows SAM has no semantic awareness to combine these masks. The narrow gap between the top and middle needs no segment. (Here, SAM+ correctly avoids segmenting it.) For the top layer (Pt / C), SAM generates masks for several subregions inside a larger region. The result implies that a single (atomically non-composite) material may contain a second material, which is semantic nonsense. This behavior is called ‘ambiguity’. Ambiguity arises when a SAM prompt causes uncertainty about how to treat a region and its immediate surrounding (parent region). In response, SAM generates masks for both the region and its parent.
@@ 4.1 Data and model
We acquired labels for 82 scanning TEM images, (mostly 1024 1024) but some with (2048 2048) pixel resolution, from two domain experts (Spurgeon, 2024). Each image contains 2 - 4 materials from among compounds including La:SrTiO, SrTiO, Ge, Nb:SrTiO, WO, LaFeO, LaMnO, LaSrFeO, and Pt / C captured in annular dark-field (ADF) and high-angle annular dark-field (HAADF) modes. Moreover, some images show vacuum regions while some material surfaces show certain defects. As shown in Fig. 3, the experts divided the images into 3 difficulty classes based on the intricacy of the texture patterns.
@@ 4.3 Results
Upon closer inspection of images that are hard to segment, we discovered why SAM and SAM-I-Am were not performing well. Certain images had very narrow ground truth masks that made it difficult to detect given our sparse grid prompt. Others appear virtually indistinguishable. In Fig. 7, for instance, the La:SrTiO and SrTiO crystalline surfaces seem identical for the untrained eye. Moreover, in some cases, unidentified defects that appear in the middle of other crystalline surfaces happen to evade SAM’s radar.
@@ 5.2 Semantic booster with supervised classifier
We experimented with an alternative booster that utilizes a supervised crop classifier in lieu of the unsupervised crop clustering approach in Fig. 2. The classifier was trained on simulated crops of SrTiO and Ge crystalline structures and applied as crop classifier preceding the mask-merging step.

### v2

@@ 1 Introduction
Unfortunately, it has been challenging to develop domain-specific analytics that extract meaningful information from raw microscope data (Kalinin et al., 2023b). A central problem in TEM analysis is the accurate and precise segmentation of microstructural image features (Akers et al., 2021; Stuckner et al., 2022). This task is difficult to automate because of the millions of potential crystal structures, defects, and artifacts in microscopy data(Spurgeon et al., 2021). Complex interactions between the electron beam and sample (Spurgeon et al., 2021) resulting from varying imaging conditions, sample orientation, sample thickness, and detector configuration, greatly change the representation of an object in collected data. As a result, unknown images are often poorly described by (limited) prior datasets.
This data sparsity is a fundamental barrier to the use of many ML-based computer vision approaches. While low-cost synthetic images generated via approaches such as multislice simulation (Rangel DaCosta et al., 2021) can be used to train models, these require extensive a priori assumptions about samples and imaging conditions. Alternatively, few-shot approaches based limited human-generated feature labels have shown some success but can fail to generalize (Akers et al., 2021). Thus, there is a critical need for new analytics to accelerate material discovery in TEM, particularly in emerging automated microscopy that combines analytics with theory-driven hypothesis generation (Liu et al., 2022).
Here we introduce an approach that harnesses the power of a zero-shot foundation model for TEM image segmentation. Most state-of-the-art TEM-segmentation techniques use convolutional neural architectures such as U-Net (Ronneberger et al., 2015). Models are trained on microscopy data that is carefully labeled by domain experts with substantial time and cost. Moreover, it remains challenging for models trained in such manner to adapt to out-of-distribution data.
@@ 1.1 Related work
(Akers et al., 2021) proposes a semi-supervised few-shot machine learning approach for TEM segmentation of three oxide material systems, (1) SrTiO3 / Ge epitaxial heterostructures, (2) La0.8Sr0.2FeO3 thin films, and (3) MoO3 nanoparticles. Given a fine-tuning set of less than 10 tiny crops (chips) of the image that represent micro-structural features, their model is able to make mask predictions comparable to those of domain experts. However, such masks are limited to chip-level representations. (Stuckner et al., 2022), proposes transfer learning from MicroNet (a large dataset of microscopy images) for improving the mean intersection over union (mIoU) of microscopy segmentation tasks. They show that models pre-trained on their dataset exhibited superior generalization to new microscopy images compared to models trained on other datasets like ImageNet.
@@ 1.2 Overview
2 Results and Discussion
@@ 2.1 Limitations of SAM+
SAM+ provides a mechanism to specify parameters that control autoprompts. Behind the scenes, it dissects an input image into layers where each layer contains equal sized crops of the image obtained by applying a sliding crop window. Next, each crop belonging to layeri is prompted by projecting (overlaying) an grid of points to identify potential regions. A typical parameter, like the ‘prompt grid length’, for instance, specifies the size of input prompt for the first layer (). Given prompt density for the first layer’s crop, crops of subsequent layers will be prompted with sparser points. i.e. .
Unfortunately, SAM+’s vanilla parameters are inadequate for achieving microstructure segmentation. Consider Fig. 1, which shows a sample with three distinct materials stacked horizontally. SAM+’s segmentation is on the left. For the middle layer (SrTiO3), SAM+ generates two separate masks. However, the whole middle section is actually the same material ‘split’ by an antiphase boundary defect (typically of special interest). This shows SAM has no semantic awareness to combine these masks. The narrow gap between the top and middle needs no segment. (Here, SAM+ correctly avoids segmenting it.) For the top layer (Pt / C), SAM generates masks for several subregions inside a larger region. The result implies that a single (atomically non-composite) material may contain a second material, which is semantic nonsense. This behavior is called ‘ambiguity’. Ambiguity arises when a SAM prompt causes uncertainty about how to treat a region and its immediate surrounding (parent region). In response, SAM generates masks for both the region and its parent.
@@ 3.1 Data and model
We acquired labels for 82 scanning TEM images, (mostly 1024 1024) but some with (2048 2048) pixel resolution, from two domain experts (Spurgeon, 2024). Each image contains 2 - 4 materials from among compounds including La:SrTiO3, SrTiO3, Ge, Nb:SrTiO3, WO3, LaFeO3, LaMnO3, La0.8Sr0.2FeO3, and Pt / C captured in annular dark-field (ADF) and high-angle annular dark-field (HAADF) modes. Moreover, some images show vacuum regions while some material surfaces show certain defects. As shown in Fig. 3, the experts divided the images into 3 difficulty classes based on the intricacy of the texture patterns.
@@ 3.3 Results
Upon closer inspection of images that are hard to segment, we discovered why SAM and SAM-I-Am were not performing well. Certain images had very narrow ground truth masks that made it difficult to detect given our sparse grid prompt. Others appear virtually indistinguishable. In Fig. 7, for instance, the La:SrTiO3 and SrTiO3 crystalline surfaces seem identical for the untrained eye. Moreover, in some cases, unidentified defects that appear in the middle of other crystalline surfaces happen to evade SAM’s radar.
@@ 4.2 Semantic booster with supervised classifier
We experimented with an alternative booster that utilizes a supervised crop classifier in lieu of the unsupervised crop clustering approach in Fig. 2. The classifier was trained on simulated crops of SrTiO3 and Ge crystalline structures and applied as crop classifier preceding the mask-merging step.